OpenStudy (anonymous):

solve this nonlinear differential equation: y''+(y'^2)+1=0 please show your work

myininaya (myininaya):

so y' is squared?

OpenStudy (anonymous):

yes it is

myininaya (myininaya):

so its (y')^2

OpenStudy (anonymous):

yes thats right

myininaya (myininaya):

ok let me think

OpenStudy (anonymous):

al right no problem but it would be helpful if you have had a course in differential equations otherwise it will be really challenging

OpenStudy (anonymous):

here is a hint...this involves using u=y' to reduce it to first order equation but it's tricky

myininaya (myininaya):

i think i got it

myininaya (myininaya):

one sec let me look at it to make sure

OpenStudy (anonymous):

ok no problem take your time you will get medal for this...promise

myininaya (myininaya):

i kind of want to type what i did but its wrong maybe someone can say where i went wrong

OpenStudy (anonymous):

i know how to work the problem...i have it done but don't know if its right because my answer does not match with the solution of this problem...we can help each other help don't worry just post what you have..

OpenStudy (anonymous):

dy/dx=v dv/dy=F(y,v) d^2y/dx^2=dv/dx=dv/dy(dy/dx)=vdv/dy

OpenStudy (anonymous):

i just need to know what you got for dy/dx ?

OpenStudy (anonymous):

you need to do Vdv/dy=-V^2-1

OpenStudy (anonymous):

i get dy/dx=(e^(2c1-2y)-1)^(1/2)

OpenStudy (anonymous):

V/V^2-1dv=dy

OpenStudy (anonymous):

noxi i am sorry i kind of do not understand what you are saying...are you differentiating?

OpenStudy (anonymous):

1/2ln(V^2-1)=y+C

OpenStudy (anonymous):

ok sorry i remember doing this last year its a special case when you only have one variable

OpenStudy (anonymous):

i get 1/2ln(v^2+1)=-y+c

OpenStudy (anonymous):

its called second order equations with independent variable missing

OpenStudy (anonymous):

my u is ur v...

OpenStudy (anonymous):

ur on the right track noxi but seems like you missed a sign somewhere but keep going and post the solution y now

OpenStudy (anonymous):

also let me know what you get for your dy/dx before you solve for y

OpenStudy (anonymous):

lets see 1/2ln(V^2-1)=-y+C ln(v^2-1)=-2(y)+C (v^2-1)=e^(-2y+C) v^2=e^(-2y+C)-1 dy/dx=SQRT(e^(-2y+C)-1)

OpenStudy (anonymous):

so yeah i got the same thing

OpenStudy (anonymous):

seems good so far...just solve for y

OpenStudy (anonymous):

aya are you there?

OpenStudy (anonymous):

yes dy/dx=SQRT(e^(-2y+C)-1) \[dy/\sqrt(e^(-2y+C)-1)=dx\] \[dy/\sqrt(C1e^(-2y0-1)=dx\] sin^-1(e^-y)=x?

OpenStudy (anonymous):

no no no no never mind forget that last part that was stupid

OpenStudy (anonymous):

shoshny and mandolino what are you guys talking about?

OpenStudy (anonymous):

aya can you post whatever you have so far?

OpenStudy (anonymous):

MeEngr what are you gettinG?

OpenStudy (anonymous):

shoshny i don't know what you talking about man

OpenStudy (anonymous):

this looks like an ugly integral \[\int\limits_{}^{}dy/\sqrt(e^(-2y)-1)\]

OpenStudy (anonymous):

yes it does very ugly indeed...i cant even type here here but it has natural log function in it

OpenStudy (anonymous):

=x i would leave it like that lol

OpenStudy (anonymous):

i typed it in wolfram and theres imaginarys in there

OpenStudy (anonymous):

yea true but you noxi whats weird is that if you solve this nonlinear differential equation on wolfram http://www.wolframalpha.com/input/?i=y%27%27%2B%28y%27^2%29%2B1%3D0 you get a really simple answer but it does not match with what we got so i am confused...is wolfram right?

OpenStudy (anonymous):

Vdv/dy=-V^2-1 do we agree up to heRE/

OpenStudy (anonymous):

i agree with u all the way up to dy/dx we are right i believe so whatever we do after that is also right but why is the solution from wolfram so different its really weird..

OpenStudy (anonymous):

did you look at the solution in wolfram?

OpenStudy (anonymous):

aya can u please post what you have so far?

OpenStudy (anonymous):

dv/dy=-(v+(1/v)) dv/(v+1/v)=-dy dy=-dv/(v^2+1/V) dy=-vdv/(v^2+1) y=-1/2ln(V^2+1) e^2y=1/v^2+1 1/e^2y-1=v^2 (1-e^2y)/e^2y=v^2 sqrt(1-e^2y)/e^2y=v=dy/dx

OpenStudy (anonymous):

noxi...what did you substitute for y' and y''?

OpenStudy (anonymous):

i think dx=sqrt(e^2y)dy/sqrt(e^2y-1) now integrate and wolfram alpha gave a do able integral check me on my math to make sure thats correct is http://www.wolframalpha.com/input/?i=sqrt%28e%5E2y%29%2Fsqrt%28e%5E2y-1%29

OpenStudy (anonymous):

i think something got screwed up in your calculations...

OpenStudy (anonymous):

isnt the prob is like this? yy''+(y'^2)+1=0 here the answer is equation of circle (x- 1/2)^2+(y+ 1/4)^2 =1/16 answer

OpenStudy (anonymous):

noxi answer my previous question

OpenStudy (anonymous):

i subbed y'=dy/dx=V y''=dv/dx=dv/dy(dy/dx)=Vdv/dy

OpenStudy (anonymous):

thats correct noxi...but your dy/dx looks fishy

OpenStudy (anonymous):

which one? when i did my calculations

OpenStudy (anonymous):

i think we did this correctly and the integral is horrible, i am off to bed, good luck and sorry i couldnt get the integral right

OpenStudy (anonymous):

y+c=-1/2ln(V^2+1) -2y-2c=ln(V^2+1) e^(-2y-2c)=sqrt(V^2+1)

OpenStudy (anonymous):

you had a problem with constant of integration and sign...but thanks for help appreciate it

OpenStudy (anonymous):

Ce^(-2y)=e^(-2y-2c)

OpenStudy (anonymous):

your sqrt(is in the wrong PLace)

OpenStudy (anonymous):

yea but here we can not bring the constant in front of e because of other problems involving boundary conditions it does not stay consistent but other than that and a few sign problems your right

OpenStudy (anonymous):

i wanted to confirm i am headed the right direction and you made me feel that i am working this right...so thanks for showing me that..

OpenStudy (anonymous):

hey aya can you send me a message in group chat...i am just wondering if you have any solution yet

OpenStudy (anonymous):

mark i am sorry about late reply but how did you arrive at that solution?

myininaya (myininaya):

no i'm still working on it i found my mistake with that one way

OpenStudy (anonymous):

ok can i see how far you have gotten...

myininaya (myininaya):

i will only show something if i have something good to show

myininaya (myininaya):

and i do not yet

OpenStudy (anonymous):

your welcome sorry i am now off to bed, good luck and problems like these on tests arent this ugly.

OpenStudy (anonymous):

isnt the prob is like yy'+(y')^2+1=0 yy'+(y')^2=-1 MULT BY 2 2[yy'+(y')^2]=-2 2+2[yy'+(y')^2]=0 2+0+2[yy'+(y')^2]=0 integrating 2x-1+0+2yy'=0 integrating x^2-x+ 1/4 +y^2= 1/16 (x- 1/2)^2+(y+ 1/4)^2 =1/16 answer

OpenStudy (anonymous):

sory m pc kept hanging up on me..lol

OpenStudy (anonymous):

ME hope thats what you are looking for....good luck now

myininaya (myininaya):

mark what happened to y''?

OpenStudy (jamesj):

This is pretty straight forward by substituting u = y' Then \[u' + u^2 + 1 = 0 \] which is a separable equation in u. Solve for u then integrate once to obtain y. Or am I missing something?

OpenStudy (anonymous):

no you need to sub y'' as well

OpenStudy (anonymous):

if u=dy/dx du/dx=du/dy(dy/dx)=udu/dy

OpenStudy (jamesj):

If u = y', then u' = y'' and the equation is \[u' + u^2 + 1 = 0\] Separating variables \[du/(u^2+1) = -dx\] and integrating \[\tan^{-1} u = - x + c_1 \ \hbox{ namely } \ u(x) = \tan(-x+c_1)\] Now, \[dy/dx = u \ \hbox{ hence } y = \int\limits \tan(-x+c_1) \ dx = \ln (\cos (-x+c_1)) + c_2\] It is easy to verify this is a solution to the equation: \[y' = \tan (-x+c_1) \ \hbox{ and} \ y'' = -\sec ^2(-x+c_1)\] hence \[y'' + (y')^2 + 1 = -\sec ^2(-x + c_1) + \tan^2(-x + c_1) + 1 = 0\] as\[\tan^2 \theta + 1 = \sec ^2 \theta \ , \ \ \hbox{for all } \ \theta\]

myininaya (myininaya):

oh wait a minute... you divided both sides by u^2+1 and so we get u'/(u^2+1)+1=0 omg brillant

myininaya (myininaya):

teach me your ways james

OpenStudy (anonymous):

Wow, that was amazing!

myininaya (myininaya):

i looked at this last night and i was trying the only way i know and was failing and then james comes in shows a freaking delicious way and i'm envious and jealous at the same time

OpenStudy (anonymous):

I was trying by substituting \(y=\frac{u'}{u}\). Not sure of that was gonna work, but I got \(u''+u=0\) which is very easy to solve and then substitute back.

OpenStudy (jamesj):

Ha, well. I'm glad I've still got it. *cracks knuckles* Next?

myininaya (myininaya):

ok i'm going to appoint james as my second zarkon

OpenStudy (anonymous):

I fanned you James! Not many people get that :P

myininaya (myininaya):

lol i fanned james too

OpenStudy (anonymous):

good job james, turning a nonlinear eq into a linear eq. great eye

OpenStudy (jamesj):

Not quite:\[ u' + u^2 + 1 = 0 \]is not linear, but it is separable.

OpenStudy (anonymous):

thats what i meant sep not linear lets see james no one else answered this can you find the annihilator for sin(x)e^(-x) i know itS D^2-2D+A^2+B^2 but i wasnt getting it to =0